### v1

# Feel the Difference? A Comparative Analysis of Emotional Arcs in Real and LLM-Generated CBT Sessions

Abstract
Synthetic therapy dialogues generated by large language models (LLMs) are increasingly used in mental health NLP to simulate counseling scenarios, train models, and supplement limited real-world data. However, it remains unclear whether these synthetic conversations capture the nuanced emotional dynamics of real therapy. In this work, we conduct the first comparative analysis of emotional arcs between real and LLM-generated Cognitive Behavioral Therapy dialogues. We adapt the Utterance Emotion Dynamics framework to analyze fine-grained affective trajectories across valence, arousal, and dominance dimensions. Our analysis spans both full dialogues and individual speaker roles (counselor and client), using real sessions transcribed from public videos and synthetic dialogues from the CACTUS dataset. We find that while synthetic dialogues are fluent and structurally coherent, they diverge from real conversations in key emotional properties: real sessions exhibit greater emotional variability, more emotion-laden language, and more authentic patterns of reactivity and regulation. Moreover, emotional arc similarity between real and synthetic speakers is low, especially for clients. These findings underscore the limitations of current LLM-generated therapy data and highlight the importance of emotional fidelity in mental health applications. We introduce RealCBT, a curated dataset of real CBT sessions, to support future research in this space.
Feel the Difference? A Comparative Analysis of Emotional Arcs in Real and LLM-Generated CBT Sessions
Xiaoyi Wang1, Jiwei Zhang1, Guangtao Zhang2, Honglei Guo3 1Department of Computer Science, Shantou University, Shantou, Guangdong, China 2Department of Automation, Tsinghua University, Beijing, China 3BNRIST, Tsinghua University, Beijing, China
1 Introduction
Mental disorders pose a major global health challenge, affecting nearly one in eight individuals worldwide WHO . As demand for mental health services continues to rise, a significant barrier remains: the shortage of trained counselors. To help bridge this gap, large language models (LLMs) such as ChatGPT and Gemini are increasingly explored as conversational agents capable of simulating therapeutic interactions and supporting clients in need. The effectiveness of these systems, however, hinges heavily on the quality and diversity of training data—particularly, counseling dialogues grounded in real-world counseling practice.
Unfortunately, access to real counseling dialogues remains extremely limited due to stringent privacy, ethical, and legal constraints. This scarcity of authentic data has led to the widespread adoption of synthetic therapy dialogues generated by LLMs. These synthetic interactions are now commonly used in mental health NLP applications for training, evaluation, and scenario simulation. For example, datasets like CACTUS Lee et al. (2024) incorporate psychological principles from Cognitive Behavioral Therapy (CBT) to structure model-generated conversations. While these synthetic dialogues are often fluent and structured around known therapeutic techniques, it remains unclear whether they capture the nuanced emotional dynamics that characterize real counseling sessions.
Emotion plays a central role in both diagnosis and treatment. In CBT, client emotions reveal cognitive distortions, inform intervention strategies, and signal therapeutic progress. A meaningful session often involves shifts in emotional state—such as the alleviation of distress or the emergence of insight. These emotional trajectories, or emotional arcs, offer a powerful lens through which to assess the depth, authenticity, and therapeutic quality of a dialogue. It is thus important to understand the extent to which synthetic therapy conversations mirror the emotional dynamics observed in real sessions.
Prior work has demonstrated that emotional arc analysis can yield rich insights into narrative structure across various domains, including novels Vishnubhotla et al. (2024); Ohman et al. (2024); Teodorescu and Mohammad (2023); Mohammad (2012), social media posts Vishnubhotla and Mohammad (2022), and movie scripts Hipson and Mohammad (2021). However, emotional arcs remain largely unexplored in the context of psychotherapy, and no prior study has directly compared real and synthetic counseling dialogues from an emotional dynamics perspective.
To address this gap, we apply emotion trajectory analysis to CBT conversations from two sources: real-world sessions transcribed from public video-sharing platforms (e.g., YouTube and Vimeo) and synthetic dialogues from the CACTUS dataset Lee et al. (2024). We examine emotional arcs from two perspectives: (1) The Emotion Dynamics of Real vs. Synthetic Therapy Dialogues: How does the emotion change in real and LLM-generated CBT dialogues: for entire dialogues, for counselors, and for clients? (2) Real vs. Synthetic Speakers: How do emotional trajectories differ between real and synthetic dialogues when analyzed separately for counselors and clients (i.e., real vs. synthetic counselors; real vs. synthetic clients)?
To answer these questions, we conduct the first in-depth comparison of emotional dynamics between real and LLM-generated CBT dialogues. We adapt the Utterance Emotion Dynamics (UED) framework  Teodorescu and Mohammad (2023); Hipson and Mohammad (2021) to measure emotion trajectories over time within counselor–client interactions.
Our analysis reveals that although synthetic dialogues are structurally fluent, they diverge from real sessions across several key emotional dimensions. Real therapy sessions display greater emotional variability, more emotion-laden language, and more authentic patterns of emotional reactivity and regulation. Additionally, the emotional arc similarity between real and synthetic speakers is low, with correlation scores typically near zero—indicating poor alignment in how emotions evolve throughout the conversation.
These findings suggest that while current LLM-generated dialogues may capture surface-level therapeutic cues, they fall short in reproducing the nuanced, co-constructed emotional flow that characterizes real-world therapy, especially in the client role.
Our contributions are twofold: (1) An annotated dataset of real-world CBT dialogues—RealCBT [^1] We are releasing the dataset.; (2) A comparative analysis of emotion dynamics between real and LLM-generated CBT sessions.
2 Related Work
2.1 Cognitive Behavior Therapy (CBT)
Cognitive Behavioral Therapy (CBT) is a widely adopted, evidence-based form of psychotherapy used to treat a variety of psychological disorders, including depression, anxiety, and addiction Beck et al. (2011). The core of CBT aims to help clients identify and challenge maladaptive thought patterns, and subsequently restructure these cognitive distortions to promote more realistic and adaptive thinking, ultimately facilitating emotional and behavioral change Carroll and Kiluk (2017); Longmore and Worrell (2007). Given its structured format and demonstrated effectiveness, CBT has increasingly been adapted into virtual agents to support individuals with limited access to in-person therapy.
2.2 Synthetic Mental Health Counseling Dialogue Datasets
Training CBT-based conversational models requires high-quality therapy dialogue datasets. However, such data are difficult to obtain due to privacy, ethical, and legal constraints. To address this limitation, recent studies have explored the use of LLMs to generate synthetic therapy dialogues that can supplement scarce real-world data. Earlier work Sharma et al. (2023); Maddela et al. (2023); Sun et al. (2021) has focused on generating single-turn counseling dialogues based on CBT or other psychological strategies. More recent research has shifted toward multi-turn generation that better simulates the interactive nature of real therapy sessions Lee et al. (2024); Xiao et al. (2024); Qiu et al. (2024). Among these, CACTUS (CBT-augmented Counseling Chat Corpus) stands out as a publicly available multi-turn dataset designed to capture the flow and depth of CBT-style conversations Lee et al. (2024). It uses LLMs to simulate counselor–client interactions guided by CBT theory and therapeutic intent. Given its theoretical grounding and multi-turn structure, we adopt CACTUS as the synthetic benchmark dataset for comparison with real-world CBT dialogues in this study. Although CACTUS has been evaluated favorably on key metrics such as helpfulness and empathy, recent work Lee et al. (2025) raises concerns that LLM-generated therapy dialogues may lack the richness and diversity of emotional expression observed in real interactions. This underscores the need to better understand how synthetic dialogues compare to real-world CBT conversations, particularly in terms of emotional dynamics, to guide the design of more realistic and effective counseling models.
2.3 Lexicon-based Emotion Dynamics Computation
Emotion dynamics is a psychological framework for measuring how an individual’s emotional state evolves over time Vishnubhotla et al. (2024); Kuppens and Verduyn (2017); Hollenstein (2015). In NLP, lexicon-based approaches are among the most widely used techniques for modeling emotion trajectories in spoken or written language. Prominent resources include LIWC Tausczik and Pennebaker (2010), WordNet-Affect Bobicev et al. (2010), SentiWordNet Baccianella et al. (2010), VADER Hutto and Gilbert (2014), and the NRC Emotion Lexicons Mohammad (2018). These lexicons provide emotion-related scores at the word level, enabling the quantification of emotional trends across text.
Lexicon-based methods have been validated in various contexts and shown to achieve high reliability in capturing emotional dynamics Ohman et al. (2024); Teodorescu and Mohammad (2023); Öhman (2021). In this study, we adopt a lexicon-based approach, specifically using the NRC Valence, Arousal, and Dominance (VAD) Lexicon Mohammad (2025, 2018) in conjunction with the Utterance Emotion Dynamics (UED) framework Hipson and Mohammad (2021). This combination allows us to compute the emotional valence, arousal, and dominance at the utterance level, and to model how these emotional states change over the course of each counseling session.
Previous research has successfully applied similar methods to track emotional arcs in literary novels Vishnubhotla et al. (2024); Teodorescu and Mohammad (2023); Mohammad (2012), social media narratives Vishnubhotla and Mohammad (2022), and movie scripts Hipson and Mohammad (2021). We extend this line of work by applying NRC and UED metrics to therapeutic dialogues, allowing us to compare emotional progression in real and synthetic CBT conversations.
3 Real-world CBT Dialogue Dataset
3.1 Data Collection
Due to privacy, ethical, and legal constraints, publicly available CBT dialogue datasets are extremely limited. To enable comparison with LLM-generated data, we collected real-world CBT Dialogues by identifying video clips from public video-sharing platforms such as YouTube and Vimeo. We restricted our search to videos explicitly labeled as CBT-based counseling sessions. For a detailed description of the selection criteria, please see Appendix A.1. In total, we collected 76 video-based CBT dialogues that met these criteria and we name this dataset as RealCBT. Detailed summary statistics of RealCBT is provided in Table 1 in Appendix A.
3.2 Preprocessing and Transcription
All videos were first converted into a standard MP4 format and then preprocessed to remove non-conversational, user-generated content such as introductory titles, animations, and narration. We used the Transkriptor [^2] https://transkriptor.com/ service to generate initial transcripts for each video. To ensure transcript quality and temporal alignment, each transcript was manually reviewed and corrected to accurately reflect the spoken dialogue in the video. Additionally, we reviewed metadata associated with each video—such as the number of views, likes, and information about the channel owner—to verify that the content originated from credible sources. These indicators collectively suggest that the selected videos represent professionally conducted CBT sessions.
3.3 Metadata Annotation
Similar to CACTUS, we annotated each dialogue with the following attributes from video metadata: (1) client problem, (2) client gender, and (3) overall client attitude toward the session. We employed three state-of-the-art language models—ChatGPT-4o Mini, Grok-v3, and Gemini 2.0 Flash—to independently infer these tags for each dialogue. A majority voting scheme was then used to determine the final label for each attribute. To evaluate the accuracy of this automated annotation approach, we manually reviewed a randomly selected subset comprising 30% of the dialogues. The automated predictions were fully consistent with the human annotations in this subset, yielding an estimated labeling accuracy of 100%. Detailed distribution of the metadata of RealCBT is provided in Table 1 in Appendix A.
4 Emotion Dynamics Computation
We adapt emotion dynamics methods from prior work by Mohammad and colleagues Vishnubhotla et al. (2024); Teodorescu and Mohammad (2023); Hipson and Mohammad (2021), who demonstrated robust approaches for capturing emotional arcs in narratives and movies. Specifically, we adopt the Utterance Emotion Dynamics (UED) framework to compute emotion metrics from sequences of utterances, treating each role (counselor or client) as a separate character trajectory.
Given that CBT sessions are considerably shorter than novels or long-form narratives, we adjust the emotion arc computation accordingly. Following prior work Hipson and Mohammad (2021); Vishnubhotla et al. (2024), we apply a rolling window of 10 words, advancing one word at a time, to compute fine-grained emotion values across each dialogue. At each windowed step, we estimate the emotional state using version 2 of the NRC Valence, Arousal, and Dominance (VAD) Lexicon Mohammad (2025), which provides word-level emotion scores for the three affective dimensions: valence (V), arousal (A), and dominance (D). Each V/A/D score ranges from –1 to 1, where 1 indicates a strong positive association and –1 indicates a strong negative association with the corresponding affective dimension.
From the resulting VAD time series, we derive a set of Utterance Emotion Dynamics (UED) metrics Hipson and Mohammad (2021) that capture the shape and character of emotion progression in therapy dialogues: (1) Emotion Mean: The average V/A/D score across a dialogue or speaker’s arc; (2) Emotion Variability: The standard deviation of V/A/D scores, representing emotional fluctuation or richness; (3) Average Displacement Length: Measures the average number of emotion words spoken more or less than usual. Longer displacements suggest stronger or more persistent emotional shifts; (4) Emotion Rise Rate: Captures how quickly and intensely a speaker shifts into an emotional state (emotional reactivity or sensitivity); (5)Emotion Recovery Rate: Indicates how quickly a speaker returns to their baseline emotional state after a shift (emotion regulation).
Based on the UED metrics and NRC VAD Lexicon, we compute three types of emotional trajectories per CBT session: (1) the overall dialogue-level arc, (2) the counselor’s arc, and (3) the client’s arc. These arcs are computed separately for both real and synthetic CBT sessions, allowing us to compare matched pairs across datasets (i.e., real vs. synthetic dialogues, real vs. synthetic counselors, and real vs. synthetic clients).
5 Experimental Setup
Figure 1: Boxplots showing the distributions of the mean and variability for each of the three affective dimensions across three comparisons: Real vs. Synthetic Dialogues, Real vs. Synthetic Counselors, and Real vs. Synthetic Clients.
CACTUS consists of 31,564 dialogues with roughly balanced distributions across attributes. In contrast, our real-world CBT dataset is relatively small and exhibits significant imbalance, particularly in terms of client gender and overall attitude toward therapy. Among the 76 collected dialogues, 84.2% of clients are female (64 out of 76), and 90.8% express a positive attitude toward the counseling process (69 out of 76). Only 12 dialogues feature male clients, and just 2 involve clients with a negative attitude—making these subgroups too sparse to support reliable statistical comparisons.
While it is theoretically possible to match gender and attitude proportions when sampling from the synthetic dataset, the severe imbalance in the real dataset would render such comparisons statistically unreliable and potentially misleading.
To ensure a fair and meaningful comparison, we instead adopt a problem-focused sampling strategy. We focus exclusively on the top three most frequent client concerns in our dataset: anxiety and fear (25 dialogues), self-esteem and confidence issues (19), and relationship-related concerns (14). Together, these categories account for 76.3% of the dataset (58 out of 76 dialogues). For a detailed description of the summary statistics and distribution of RealCBT, please see Table A.2 and Table A.2 in Appendix A. We sample synthetic dialogues from CACTUS to match this distribution.
To enhance the robustness of our analysis, we repeat the sampling process 10 times without replacement, generating 10 non-overlapping subsets of synthetic dialogues. All quantitative results are averaged across these runs to account for sampling variability.
6 Experimental Results
6.1 The Emotion Dynamics of Real vs. Synthetic Therapy Dialogues
We analyze and compare the following UED metrics across valence, arousal, and dominance: Emotion Mean, Variability, Displacement Length, Rise and Recovery Rates. To assess statistical significance between real and synthetic groups, we apply the Mann–Whitney U test for non-parametric pairwise comparisons, using a p-value threshold of 0.05. Figure 1 presents boxplots illustrating the distributions of Emotion Mean and Variability for each role—dialogue, counselor, and client—in both real and synthetic sessions. Blue dots indicate the aggregated Emotion Mean and Variability for each affective dimension, summarizing overall emotional trends. Detailed summary statistics for Emotion Mean, Emotion Variability, and all UED metrics across VAD are provided in Table A.2 - A.2 in Appendix A.
6.1.1 Emotion Mean
As shown in Figure 1, synthetic speakers (the whole dialogue, counselor, and client) generally sound more emotionally elevated or intense than real human speakers. An exception is observed in valence, where real and synthetic clients show similar scores. We use the Mann–Whitney U test to perform pairwise comparisons between real and synthetic groups on Emotion Mean, across 10 independently sampled synthetic datasets. The test results support this observation.
For entire dialogues (as a whole speaker), we found significant differences in arousal in 90.0% of cases (median p-value << 0.001, average effect size: 0.264)
For counselors, we observed significant differences in valence in all comparisons (median p-value: 0.00019, range: $1.361\times 10^{-7}$ to 0.0035, average effect size: 0.357). For clients, we found significant differences in arousal in 70.0% of comparisons (median p-value: 0.02, range: 0.001225 to 0.09596, average effect size: 0.226) and in valence in 90.0% (median p-value: 0.0046, range: $4.178\times 10^{-7}$ to 0.0655), average effect size: 0.260).
6.1.2 Emotion Variability
For entire CBT dialogues, real sessions exhibited significantly greater variability in arousal than synthetic ones, indicating more fluctuation or richness in real speakers. All 10 Mann–Whitney U comparisons showed significant differences, with a median p-value of $7.677\times 10^{-6}$ (range: $5.053\times 10^{-8}$ to $0.007753$) and an average effect size of 0.836, indicating a strong effect. In contrast, valence and dominance variability were similar between real and synthetic dialogues, with only partial significant differences (i.e., 5/10 in dominance and 7/10 in valence).
For counselors, real speakers also showed significantly greater variability in both arousal and valence compared to synthetic counselors. All comparisons were significant. For arousal, the median p-value was $1.486\times 10^{-5}$ (range: $4.97\times 10^{-7}$ to $0.0004315$), with an average effect size of 0.406. For valence, the median p-value was $5.699\times 10^{-6}$ (range: $1.561\times 10^{-9}$ to $0.001349$), with an average effect size of 0.422. For dominance, real counselors tended to exhibit greater variability, though the differences were not consistently significant across the comparisons.
For clients, synthetic sessions tended to show greater variability in valence, while real clients showed slightly more variability in arousal. However, these differences were only partially significant (i.e., 6/10 for arousal and 1/10 for valence). Dominance variability was similar across both groups, with no significant differences observed.
6.1.3 Displacement Lengths
We found that speakers in real dialogues talk significantly more emotion words than the ones in synthetic dialogues in both arousal and dominance. In particular, for arousal, all comparisons showed significant differences, with a median p-value of 0.0001618 (range: $4.643\times 10^{-6}$ to 0.001617) and an average effect size of 0.354. For dominance, 70% of comparisons were significant, with a median p-value of 0.02345 (range: 0.0001469 to 0.1449) and an average effect size of 0.232. In contrast, both real and synthetic speakers uttered similar amount of emotion words in dominance with minimal difference—only 1 out of 10 comparisons reaching significance.
Real counselors tend to speak more emotion words than synthetic counselors. 70% of comparisons showed significant differences in both arousal and valence. The median p-values were 0.0233 (arousal) and 0.007435 (valence), with average effect sizes of 0.227 in both cases. In contrast, dominance minimal showed significant differences in only 20% of comparisons, indicating similar amount of emotion words uttered by both real and synthetic counselors.
Real clients utter more emotion words than synthetic ones in the dimensions of arousal and dominance. Arousal showed consistent differences across all comparisons, with a median p-value of 0.0003099 (range: $1.76\times 10^{-6}$ to 0.007527) and an average effect size of 0.335. Dominance showed partial significant differences in 70% of comparisons, with a median p-value of 0.0163 (range: $5.556\times 10^{-5}$ to 0.1787) and an average effect size of 0.239. In contrast, valence showed no significant differences across any comparisons.
6.1.4 Emotion Rise Rate
We observed that synthetic speakers tend to respond more quickly and intensely to emotional situations than real speakers in dominance (synthetic 0.0332 vs. real 0.0311) and valence (synthetic 0.0255 vs. real 0.0240). In particular, dominance showed relative strong significant differences in 70% of comparisons with the median p-value 0.02603 (range: 0.001542 to 0.2407, average effect size: 0.204). Similarly, valence demonstrate partial significant differences (50% of comparisons, median p-value: 0.04623, effect size: 0.185). However, real speakers seem to be slightly higher arousal scores than synthetic ones with no significant differences.
Real and synthetic counselors have very similar scores in valence (0.0328 vs. 0.0310), arousal (0.0254, 0.0258), and dominance (0.0247, 0.0254), indicating that they react to emotional situations with similar intensity and speed. There were no consistence significant differences observed.
Synthetic clients seem to be more actively respond to emotional situations than real ones in the dimensions of valence (0.0349, 0.0319) and dominance (0.0262, 0.0240). Both dimensions showed significant differences in 60% of comparisons with the median p-value valence 0.02253 (range: 0.004904 to 0.1897, average effect size: 0.197) and dominance 0.02273 (range: 0.001897 to 0.5602, average effect size: 0.191). Both real and synthetic clients reacted similarly in arousal (0.0270, 0.0262) with no significant differences.
6.1.5 Emotion Recovery Rate
Figure 2: Distribution of Spearman Correlation Scores between Real and Synthetic Speakers (Upper: Real vs. Synthetic Counselors; Lower: Real vs. Synthetic Clients)
In general, synthetic and real speakers showed similar ability to return to their baseline emotional states after a shift in valence (0.0325 vs. 0.0313) and arousal (0.0254 vs. 0.0261), with no significant differences observed. However, dominance showed partial differences: 60% of comparisons were significant, with a median p-value of 0.01965 and an average effect size of 0.217, suggesting that synthetic speakers may exert slightly more emotional control in this dimension.
For counselors, emotion regulation ability was comparable across all three affective dimensions: valence (real: 0.0324 vs. synthetic: 0.0304), arousal (0.0261 vs. 0.0260), and dominance (0.0250 vs. 0.0248). No significant differences were observed.
In contrast, synthetic clients exhibited stronger emotion regulation than real clients in both valence and dominance. Valence showed significant differences in all comparisons (median p-value: $1.698\times 10^{-5}$; average effect size: 0.401), while dominance showed partial significance in 50% of comparisons (median p-value: 0.04686; effect size: 0.168). No significant differences were found in arousal, suggesting similar regulation patterns in that dimension.
6.1.6 Discussion
Synthetic sessions are generally fluent and well-structured, yet the emotional expressions of speakers still differ in subtle but important ways from those observed in real therapy interactions. In the following, we discuss several important distinctions in the emotional dynamics of real and synthetic CBT dialogues based on our comparative analysis.
First, synthetic dialogues tend to exhibit higher overall emotion means, especially in arousal and valence. This phenomenon suggests that synthetic speakers may adopt a more emotionally elevated or expressive tone. This could be a result of LLMs overemphasizing affective expression to appear engaged or empathetic. In contrast, real counselors and clients demonstrated more restrained emotional expression, aligning with the grounded and calibrated communication style expected in professional therapy.
Second, emotion variability serves as an indicator of emotional richness in dialogue. We observed consistently higher variability in real sessions, particularly in arousal and valence. This suggests that real speakers fluctuate more in their emotional tone throughout the conversation, reflecting more authentic, responsive, and context-sensitive affective expression. In contrast, synthetic dialogues tend to be more emotionally uniform, which may result from LLMs generating emotionally consistent responses without deeply modeling the evolving emotional context.
Notably, this difference in variability remains even though the synthetic dataset is ten times larger than the real dataset (580 vs. 58 dialogues). Despite the increased diversity that might be expected from a larger sample, the real dataset still exhibits greater emotional variability. This suggests that current LLMs are struggling to replicate emotion dynamics in real therapy sessions.
Third, displacement length analysis further supports this observation: real speakers used emotion-laden language more dynamically, especially in arousal and dominance dimensions. This suggests that real speakers vary their emotional emphasis in different parts of the session, whereas synthetic speakers may be more emotionally formulaic or flat in structure.
Fourth, in terms of emotional reactivity, synthetic speakers, particularly clients, exhibited faster and more intense shifts in emotion in dominance and valence. While this may appear to reflect emotional engagement, it might also signal exaggerated or less nuanced affective responses, possibly resulting from LLMs optimizing for expressive content over authentic regulation.
Relatedly, emotion recovery rate, reflecting emotion regulation, was largely comparable across real and synthetic sessions, though synthetic clients demonstrated stronger emotion regulation in valence and dominance. This pattern might again reflect stylized generation patterns that smooth or suppress emotional volatility, possibly at the cost of naturalistic variation.
Taken together, these findings suggest that synthetic dialogues approximate but do not fully replicate the emotional dynamics of real therapy conversations. Synthetic counselors show relatively close alignment with real counselors across most metrics, indicating that LLMs are reasonably good at simulating professional emotional tone. In contrast, synthetic clients diverge more notably, particularly in reactivity and regulation, which may affect the authenticity of the simulated therapeutic process.
6.2 Similarity between Real and Synthetic Speakers
In this section, we investigate how closely LLM-generated speakers—counselors and clients—mirror the emotional trajectories of real individuals by analyzing emotion arc similarity. Following the approach by Vishnubhotla et al. Vishnubhotla et al. (2024), we temporally align real and synthetic emotion arcs and compute Spearman correlation between each real and synthetic pair (i.e., real vs. synthetic counselors, real vs. synthetic clients). These correlation scores provide a measure of trajectory alignment: higher values indicate stronger similarity in emotional progression.
[…]
6.2.1 Discussion
The correlation analysis reveals that LLM-generated counselors and clients do not closely replicate the emotional trajectories of their real counterparts. Across both roles and all three VAD dimensions, the emotion arc similarity remains low, with most correlation values near zero and only a few outliers showing moderate alignment—often in the negative direction. This suggests that while LLMs may produce locally appropriate emotional expressions, they lack the nuanced emotional progression found in real therapeutic interactions. The weak alignment implies that current synthetic dialogues may fail to capture the dynamic, interpersonal emotional flow that characterizes real therapy, especially the moment-to-moment adaptation that occurs between counselor and client. Improving emotional trajectory modeling—particularly over longer conversational contexts—remains a key challenge for enhancing the realism and therapeutic value of LLM-generated dialogues.
7 Conclusion
In this work, we conduct the first systematic comparison of emotional dynamics between real and LLM-generated CBT dialogues. Using the UED framework, we analyzed emotion mean and variability, reactivity, regulation, and arc similarity across counselors and clients. Our findings show that while synthetic dialogues are fluent and structured, they diverge from real therapy in key emotional dimensions—particularly in variability and emotional arc alignment. These results highlight the limitations of current LLMs in replicating the nuanced emotional flow of real therapeutic interactions, especially on the client side. This work introduces the RealCBT dataset alongside an in-depth analysis for evaluating affective realism in synthetic therapy, paving the way for future research on more emotionally grounded and psychologically credible dialogue generation.
8 Limitations
Our study is limited by the size and distribution of the real-world CBT dataset. Despite careful curation, the dataset comprises only 76 dialogues due to the difficulty of obtaining publicly available, high-quality therapy sessions. This limited sample size constrains the generalizability of our findings. Furthermore, the dataset is highly imbalanced in terms of speaker attributes: 84.2% of the clients are female, and 90.8% exhibit a positive overall attitude toward the session. These imbalances make it challenging to conduct reliable subgroup analyses (e.g., by gender or attitude) and may introduce bias into comparative evaluations. While we mitigate some of these limitations by focusing on the top three most frequent client problem types and applying stratified sampling in our comparisons with synthetic data, the representativeness of the dataset remains a constraint. Future work should explore larger and more demographically diverse datasets to validate and extend our findings.
Appendix A Appendix
A.1 RealCBT Selection Criteria
We used keywords such as CBT counseling, CBT case study, CBT role play, CBT session, and using CBT to guide the search. To ensure that the selected videos accurately portrayed CBT-style therapy conversations, we applied the following inclusion criteria: (1) the video must feature exactly two participants—a counselor and a client; (2) the video should contain little to no narration or background music that could interfere with the dialogue; and (3) the session should focus on a behavioral or emotional issue consistent with CBT practice, such as depression, anxiety, or smoking cessation; (4) the therapy dialogue should last at least three minutes to ensure sufficient conversational context for analysis.
A.2 Three Representative Examples of Emotion Arc Correlation between Real and Synthetic Clients
Figure 3 illustrates three representative examples of emotion arc correlation between real and synthetic clients: (a) high positive correlation (strong alignment), (b) near-zero correlation (no alignment), and (c) high negative correlation (strong misalignment).
These examples demonstrate the variability in arc similarity and emphasize that strong alignment is rare in synthetic client behavior.
Figure 3: Emotion arcs of valence, arousal, and dominance for a client in three representative cases: (a) highest correlation, (b) near-zero correlation, and (c) lowest (negative) correlation between real and synthetic trajectories.
Category Subcategory Proportion (%) Count
Client’s Problem Anxiety and fear 32.89 25
Self-esteem and confidence issues 25.00 19
Relationships (romantic, family, friendships) 18.42 14
Career and work-related concerns 9.21 7
Health-related worries 5.26 4
Academic and educational concerns 3.95 3
Financial concerns 2.63 2
Health-related Worries 1.32 1
Other miscellaneous concerns 1.32 1
Client’s Attitude Positive 90.79 69
Neutral 6.58 5
Negative 2.63 2
Client’s Gender Female 84.21 64
Male 15.79 12
Total 100.00 76
Table 1: Distribution of the RealCBT dataset
Number of sessions 58 Total client words count 59,580 Total counselor words count 82,618 Total word count 142,198 Average word count per session 2,461 Total Duration (min) 907.18 Avg. Duration (min) 15.64 Table 2: Descriptive statistics of RealCBT dataset (Top 3 Client Problems)
Category Subcategory Proportion (%) Count Client’s Problem Anxiety and fear 43.10 25 Self-esteem and confidence issues 32.76 19 Relationships (romantic, family, friendships) 24.14 14 Client’s Attitude Positive 93.10 54 Negative 3.45 2 Neutral 3.45 2 Client’s Gender Female 87.93 51 Male 12.07 7 Total 100.00 58 Table 3: Distribution of Top 3 Client Problems in the RealCBT dataset
Speaker Valence Arousal Dominance Mean Var. Mean Var. Mean Var. Real Dialogue 0.2334 0.1448 -0.0907 0.1161 0.0509 0.1131 Synthetic Dialogue 0.2761 0.1480 -0.0247 0.1037 0.1130 0.1154 Real Counselor 0.2302 0.1466 -0.0804 0.1148 0.0614 0.1120 Synthetic Counselor 0.3067 0.1287 -0.0240 0.0998 0.1550 0.1097 Real Client 0.2344 0.1312 -0.1006 0.1109 0.0346 0.1052 Synthetic Client 0.2339 0.1518 -0.0257 0.1012 0.0655 0.1043 Table 4: Emotion Mean and Variability in Valence, Arousal, and Dominance across RealCBT and CACTUS
metric//sperker type Real Dialogue Synthetic Dialogue Real Counselor Synthetic Counselor Real Client Synthetic Client emo_mean 0.2334 0.2761 0.2302 0.3067 0.2344 0.2339 emo_std 0.1448 0.1480 0.1466 0.1287 0.1312 0.1518 emo_avg_peak_dist 0.0762 0.0779 0.0792 0.0695 0.0718 0.0818 emo_avg_disp_length 4.2100 4.2282 4.2996 3.9746 3.9735 4.0276 emo_rise_rate 0.0311 0.0332 0.0328 0.0310 0.0319 0.0349 emo_recovery_rate 0.0313 0.0325 0.0324 0.0304 0.0304 0.0369 emo_low_peak_dist 0.0896 0.0887 0.0914 0.0816 0.0838 0.0920 emo_low_disp_length 4.3939 4.2972 4.4628 4.3451 4.3002 4.0820 emo_low_rise_rate 0.0346 0.0356 0.0352 0.0321 0.0339 0.0381 emo_low_recovery_rate 0.0338 0.0352 0.0343 0.0328 0.0315 0.0379 emo_high_peak_dist 0.0666 0.0705 0.0705 0.0637 0.0664 0.0810 emo_high_disp_length 4.2266 4.3962 4.3971 4.0169 4.0794 4.6133 emo_high_rise_rate 0.0280 0.0312 0.0303 0.0304 0.0299 0.0321 emo_high_recovery_rate 0.0292 0.0301 0.0305 0.0285 0.0297 0.0367 Table 5: UED Metrics in Valence across RealCBT and CACTUS
metric//sperker type Real Dialogue Synthetic Dialogue Real Counselor Synthetic Counselor Real Client Synthetic Client emo_mean -0.0907 -0.0247 -0.0804 -0.0240 -0.1006 -0.0257 emo_std 0.1161 0.1037 0.1148 0.0998 0.1109 0.1012 emo_avg_peak_dist 0.0618 0.0557 0.0590 0.0555 0.0616 0.0564 emo_avg_disp_length 4.1024 3.6634 3.8828 3.5889 4.1044 3.5794 emo_rise_rate 0.0260 0.0252 0.0254 0.0258 0.0270 0.0262 emo_recovery_rate 0.0261 0.0254 0.0261 0.0260 0.0259 0.0262 emo_low_peak_dist 0.0526 0.0531 0.0516 0.0536 0.0542 0.0546 emo_low_disp_length 4.1279 3.7331 3.8291 3.7313 4.3172 3.6903 emo_low_rise_rate 0.0232 0.0244 0.0236 0.0249 0.0240 0.0248 emo_low_recovery_rate 0.0231 0.0243 0.0240 0.0248 0.0248 0.0253 emo_high_peak_dist 0.0732 0.0605 0.0723 0.0617 0.0727 0.0634 emo_high_disp_length 4.2461 3.7695 4.4102 3.8029 4.2717 3.8465 emo_high_rise_rate 0.0290 0.0263 0.0282 0.0270 0.0302 0.0278 emo_high_recovery_rate 0.0293 0.0268 0.0282 0.0274 0.0275 0.0273 Table 6: UED Metrics in Arousal across RealCBT and CACTUS
metric//sperker type Real Dialogue Synthetic Dialogue Real Counselor Synthetic Counselor Real Client Synthetic Client emo_mean 0.0509 0.1130 0.0614 0.1550 0.0346 0.0655 emo_std 0.1131 0.1154 0.1120 0.1097 0.1052 0.1043 emo_avg_peak_dist 0.0608 0.0612 0.0595 0.0592 0.0588 0.0579 emo_avg_disp_length 4.3631 4.0982 4.1802 4.0086 4.2093 3.7704 emo_rise_rate 0.0240 0.0255 0.0247 0.0254 0.0240 0.0262 emo_recovery_rate 0.0239 0.0256 0.0250 0.0248 0.0247 0.0263 emo_low_peak_dist 0.0619 0.0645 0.0614 0.0655 0.0610 0.0619 emo_low_disp_length 4.3842 4.2835 4.2933 4.3020 4.3629 4.0320 emo_low_rise_rate 0.0247 0.0260 0.0243 0.0268 0.0237 0.0269 emo_low_recovery_rate 0.0240 0.0260 0.0247 0.0252 0.0244 0.0262 emo_high_peak_dist 0.0628 0.0609 0.0609 0.0587 0.0605 0.0604 emo_high_disp_length 4.5743 4.1729 4.4043 4.2198 4.4529 4.0397 emo_high_rise_rate 0.0237 0.0250 0.0250 0.0242 0.0242 0.0261 emo_high_recovery_rate 0.0241 0.0254 0.0255 0.0247 0.0251 0.0266 Table 7: UED Metrics in Dominance across RealCBT and CACTUS